OpenStudy (anonymous):

simplify 1-sin^2x/sinxcsc x

OpenStudy (anonymous):

IDENTITIES!!! did this last week in calculus.

OpenStudy (anonymous):

does it help to know that \(1-\sin^2(x)=\cos^2(x)\) ?

OpenStudy (anonymous):

also need to know that cosecant is the reciprocal of sine so \(\sin(x)\csc(x)=1\)

OpenStudy (anonymous):

I typed the equation wrong, Its 1-sin^2x/ sinx -cscx

OpenStudy (anonymous):

lol ok lets start again

OpenStudy (anonymous):

the numerator is still \(\cos^2(x)\)

OpenStudy (anonymous):

the denominator is \[\sin(x)-\frac{1}{\sin(x)}=\frac{\sin^2(x)-1}{\sin(x)}\]

OpenStudy (anonymous):

would you multiply the numerator by the reciprocal?

OpenStudy (anonymous):

you good from there, or you want to walk through the algebra

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and when you do that, you will have a term that looks like \[\frac{1-\sin^2(x)}{\sin^2(x)-1}\] which is just \(-1\)

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

I know the answer is supposed to be -sinx, but I got cos^2 x

OpenStudy (anonymous):

i misled you about changing \(1-\sin^2(x)\) to \(\cos^2(x)\) i mean it is true but not necessary

OpenStudy (anonymous):

\[\frac{(1-\sin^2(x))\times \sin(x)}{\sin^2(x)-1}=-\sin(x)\] because \[\frac{1-\sin^2(x)}{\sin^2(x)-1}=-1\]

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

I started over, and I'm stuck on sin^2x-1 * 1-sin^2x

OpenStudy (anonymous):

ahh because you have to multiply the the reciprocal!

OpenStudy (anonymous):

\[\frac{1-\sin^2(x)}{\frac{\sin^2(x)-1}{\sin(x)}}=1-\sin^2(x)\times \frac{\sin(x)}{\sin^2(x)-1}\]

OpenStudy (anonymous):

you gotta flip the bottom one that is why it is \[\frac{(1-\sin^2(x))\sin(x)}{\sin^2(x)-1}\]

OpenStudy (anonymous):

This is where I get really confused, I have \[1-\sin ^{2}x * \frac{ sinx }{ \sin ^{2}x-1 }\]

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

wouldn't I need to multiply, so i have a common denominator? \[\frac{ 1-\sin ^{2}x }{ 1-\sin ^{2}x } * \frac{ 1-\sin ^{2} x}{ 1 }\]

OpenStudy (anonymous):

multiply means multiply in the numerator, not in the denominator i.e. it is \[\frac{1-\sin^2(x)}{1}\times \frac{\sin(x)}{\sin^2(x)-1}\] you don't find a common denominator when you multiply fractions, that is for addition

OpenStudy (anonymous):

suppose you had numbers, like \[\frac{3}{\frac{2}{5}}\] that is the same as \[3\times \frac{5}{2}=\frac{15}{2}\]

OpenStudy (anonymous):

Ohh! Thats where I was going wrong. So it would be \[\frac{ 1-\sin ^{3}x }{ \sin ^{2}x-1 }\]

OpenStudy (anonymous):

you missed the parentheses when you multiplied it is \[\frac{(1-\sin^2(x))\sin(x)}{\sin^2(x)-1}\]

OpenStudy (anonymous):

don't multiply by \(\sin(x)\) but rather cancel the top and the bottom as \(-1\)

OpenStudy (anonymous):

is this step okay?\[\frac{1-\sin^2(x)}{1}\times \frac{\sin(x)}{\sin^2(x)-1}\]

OpenStudy (anonymous):

Yes, that's what I have. I don't understand how it is -1 though. The answer is supposed to be -sinx

OpenStudy (anonymous):

ok lets go slow we got that part, now lets make it look like this \[\frac{1-\sin^2(x)}{\sin^2(x)-1}\times \frac{\sin(x)}{1}\] is that okay? i just rearranged it a bit

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ahh damn typo sorry,i meant the first part is \[\frac{1-\sin^2(x)}{\sin^2(x)-1}\]

OpenStudy (anonymous):

and that is always \(-1\) just like \[\frac{1-5}{5-1}=-1\] no matter what, that is \(-1\)

OpenStudy (anonymous):

so \[\frac{1-\sin^2(x)}{\sin^2(x)-1}\times \sin(x)=-1\times \sin(x)=-\sin(x)\]

OpenStudy (anonymous):

Okay, I got it! Thank you so much!

OpenStudy (anonymous):

whew yw now i have a question what is "46"?

OpenStudy (anonymous):

It was my entry number for competition

OpenStudy (anonymous):

hope you won

OpenStudy (anonymous):

I did, thanks.